Sphere with mass 200 g and initial velocity 2 m / s stops after 10 s under the action of friction friction force f. What is the value of this force?
1
Expert's answer
2020-12-14T07:17:21-0500

Find the acceleration:


"a=\frac{\Delta v}{t}=\frac{v-0}{t}=\frac vt."

Find the force according to Newton's second law:


"f=ma=m\frac vt=0.04\text{ N}."